ampicillin 100mg/kg/day PO divided in doses every 6 hours for a child who weighs 49 lbs 2 oz. Available is 250mg/5mL. What is the amount to administer to the child in mL.?

please show desired over have method.
Answer #1

Ampicillin 100mg/kg/ day has to be administered to the child in question

Given,

the child weighs 49 pound 2 oz,

This needs to be converted to kilograms. We know that 1 pound= 16 oz

49lbs 2 oz= (49*16) + 2= 786 oz. Since 1 oz= 35.2 kg

786 oz= 22.3 kg(approximately). For the sake of easier calculations, let us round figure the weight of the child to 22kgs

Now, we calculate the dose that needs to be administered to the child (in mg), which is,

100*22= 2200 mg/day

This dose needs to be given to the child every 6 hours,i.e. 4 times daily

hence 2200 divided by the no. of dose per day, which is 4= 550mg/ dose every 6 hours

Available form of the antibiotic is 250mg/5 mL which basically means every 5 mL of the medicine contains 250 grams of Ampicillin( 5 mL contains 50 mg of ampicillin)

Next we convert the dose from mg to mL,

550 mg / 50= 11 mL

The Child should be given 11 mL of the medicine every 6 hours.
Answer #2
11.2 mL/dose